OpenStudy (anonymous):

QUESTION?! The area of Desert A is nine times the area of Desert B. If the Sum of their areas is 4,000,000 square miles, find the area of each desert?????

OpenStudy (anonymous):

not a big answer but if the sums of the areas are 4.000.000 then let me see..

OpenStudy (anonymous):

Thank you so much! @Bella106

OpenStudy (anonymous):

what 2+2

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

think that 2.000.000+2.000.00 = 4.00.000

OpenStudy (anonymous):

so think that each sum of the areas are 2.000.000 cause if the total of the 2 areas are 4 .000.000 and your trying to find the sum for one of them ...2+2=4 like saying 2.000.000 + 2.000.000 = 4.000.000 so ur answer is 2.000.00 hope I helped!

OpenStudy (anonymous):

@Bella106 you did thank you.. But what if it's 5 times the area of B so the square mile is 1,200,000?

OpenStudy (anonymous):

lol I did 2.000.000*5 = 10 then 2.000.000 / 5 =0.4 so I think ur right